# Center-freeness of finite-step solvable groups arising from anabelian geometry

### v1

@@ 1.3. Proof of the slimness of free $m$-step solvable groups
it suffices to show that $\rho_{H}(g)=1$ (i.e., $g\in H^{[m-1]}$) for each such $H$, where $\rho_{H}:H\twoheadrightarrow H^{m-1}$ is the natural surjection. The image of $x$ in the finite quotient $\mathcal{F}^{m}/H$ has finite order. Let NN denote this order. Since $g$ commutes with $x^{n}$, it also commutes with $x^{Nn}$, and therefore $\rho_{H}(g)$ commutes with $\rho_{H}(x^{Nn})$. By the Nielsen–Schreier theorem, the inverse image $\tilde{H}$ of $H$ in $\mathcal{F}$ is again a free pro-$\Sigma$ group, and we may choose a free generating set of $\tilde{H}$ containing $x^{N}$. By A, we have $H^{m-1}\cong\tilde{H}^{m-1}$. Applying the induction hypothesis for m1m-1 to $\tilde{H}^{m-1}$ and the basis element $x^{N}\in\tilde{H}^{m-1}$, we obtain
$\mathrm{C}_{H^{m-1}}\bigl((x^{N})^{n}\bigr)=\langle x^{N}\rangle.$
On the other hand, we have $g\in(\mathcal{F}^{m})^{[m-1]}\subset(\mathcal{F}^{m})^{[2]}\subset H^{[1]}$ and hence $\rho_{H}(g)\in(H^{m-1})^{[1]}$. Note that $\langle x^{N}\rangle$ maps injectively to $(H^{m-1})^{\mathrm{ab}}$, whereas $(H^{m-1})^{[1]}$ maps trivially. Therefore,
$\rho_{H}(g)\in\langle x^{N}\rangle\cap(H^{m-1})^{[1]}=1.$
@@
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle$
@@
Finally, we consider the case $r=\infty$. Let $J$ be the directed set of finite subsets $X_{j}$ of $X$ such that $x\in X_{j}$. For each $j\in J$, let $\mathcal{F}_{j}$ be the finitely generated free pro-$\Sigma$ group on $X_{j}$ and let $\pi_{j}:\mathcal{F}\twoheadrightarrow\mathcal{F}_{j}$ be the continuous morphism sending generators in $X_{j}$ to themselves and generators in $X\setminus X_{j}$ to the identity element of $\mathcal{F}_{j}$. Additionally, let $\pi_{j}^{(m)}:\mathcal{F}^{m}\twoheadrightarrow\mathcal{F}_{j}^{m}$ be the natural projection induced from $\pi_{j}$. Then, by [12, Proposition 3.3.9], we have an isomorphism
@@
Let $g\in\mathrm{C}_{\mathcal{F}^{m}}(x^{n})$. By the finite-rank case, we obtain $\pi_{j}^{(m)}(g)\in\langle\pi^{(m)}_{j}(x)\rangle$ for all $j$. Passing to the inverse limit, we conclude that $g\in\langle x\rangle$. Thus, $\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle$ also holds when $r=\infty$. This completes the proof. ∎
For a profinite group $G$, we say that $G$ is slim if the centralizer $\mathrm{C}_{G}(H)$ of any open subgroup $H$ of $G$ is trivial. We note that slimness implies center-freeness.
@@
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$. Assume that $r\neq 1$. Then $\mathcal{F}^{m}$ is slim for all $m\in\mathbb{Z}_{\geq 2}$.
@@
$\mathrm{C}_{\mathcal{F}^{m}}(H)\subset\mathrm{C}_{\mathcal{F}^{m}}(x^{n})\cap\mathrm{C}_{\mathcal{F}^{m}}((x^{\prime})^{n^{\prime}})=\langle x\rangle\cap\langle x^{\prime}\rangle=1,$
where the last equality follows from the facts that $\langle x\rangle$ and $\langle x^{\prime}\rangle$ embed into the abelianization $\mathcal{F}^{\mathrm{ab}}$ and are distinct. This completes the proof. ∎
@@ 2. The $m$-step solvable Grothendieck conjecture
In this section, we show that the maximal $m$-step solvable quotients of the geometric étale fundamental groups of hyperbolic curves over a field of characteristic 0 are center-free (see Theorem 2.6). Moreover, we explain how to relate this result to anabelian geometry and the Grothendieck conjecture. Throughout this section, let $\Sigma$ be a non-empty set of prime numbers.
2.1. A strategy for $m$-step solvable center-freeness
@@
We record a strategy used in proofs of center-freeness for maximal $m$-step solvable quotients.
@@
    We say that $G$ is ab-faithful if, for each open subgroup $H$ of $G$ and each open normal subgroup NN of $H$, the natural morphism
    $H/N\rightarrow\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)$
@@
    Let $M$ be a (continuous) $G$-module. Then we say that the action $G\to\operatorname{Aut}(M)$ is fixed-point-free if $M^{G}=0$.
Remark $2.$
Let $G$ be a profinite group and let $m\in\mathbb{Z}_{\geq 2}$. For any open subgroup $P$ of $G$ such that $G^{[m-1]}\subset P$, let $\overline{P}\subset G^{m}$ be its image under $G\twoheadrightarrow G^{m}$. Then $P^{\mathrm{ab}}\xrightarrow{\sim}\overline{P}^{\mathrm{ab}}$ by A. In particular, the following hold:
@@
    Assume that $G$ is ab-torsion-free, and let NN be as in Definition 2.11. Then $\overline{N}^{\mathrm{ab}}$ is also torsion-free when $G^{[m-1]}\subset N$.
@@
    Assume that $G$ is ab-faithful, and let $H$ and NN be as in Definition 2.12. Assume that $G^{[m-1]}\subset N$. Then we have canonical isomorphisms
    $N^{\mathrm{ab}}\xrightarrow{\sim}\overline{N}^{\mathrm{ab}}\qquad\text{and}\qquad H/N\xrightarrow{\sim}\overline{H}/\overline{N}.$
    Hence the conjugation action of $\overline{H}/\overline{N}$ on $\overline{N}^{\mathrm{ab}}$ is also faithful.
@@
Let $G$ be a profinite group and let $m\in\mathbb{Z}_{\geq 2}$. Assume the following two conditions:
  1. (a)
    The profinite group $G$ is ab-faithful;
  2. (b)
    The conjugation action of $G^{m}$ on $(G^{m})^{[m-1]}$ is fixed-point-free.
Then $G^{m}$ is center-free.
@@
Let NN be an open normal subgroup of $G$ containing $G^{[m-1]}$, and let $\overline{N}\subset G^{m}$ be its image under $G\twoheadrightarrow G^{m}$. By the definition of the center, we have
@@
$Z(G^{m})\subset\ker\Bigl(G^{m}\to\operatorname{Aut}(\overline{N}^{\mathrm{ab}})\Bigr).$
The action of $G^{m}/\overline{N}$ on $\overline{N}^{\mathrm{ab}}$ is faithful by a (see Remark 2.22), and hence the right-hand kernel equals $\overline{N}$. By running over all such NN, we obtain
$Z(G^{m})\subset\bigcap\overline{N}=(G^{m})^{[m-1]}.$ (2.1)
On the other hand, for any $a\in G^{m}$, the condition $a\in Z(G^{m})$ is equivalent to the condition that $gag^{-1}=a$ for all $g\in G^{m}$, hence
$Z(G^{m})\cap(G^{m})^{[m-1]}=((G^{m})^{[m-1]})^{G^{m}}=1,$
where the last equality follows from b. Combining this with 2.1, we obtain $Z(G^{m})=1$. ∎
@@
Let $G$ be an ab-torsion-free profinite group. Then the conjugation action of $G^{m}$ on $(G^{m})^{[m-1]}$ is fixed-point-free for all $m\in\mathbb{Z}_{\geq 2}$.
@@
Let $\mathcal{N}$ be the set of all open normal subgroups of $G$ containing $G^{[m-1]}$. Fix $N\in\mathcal{N}$. First, we claim that the natural morphism
$(N^{\mathrm{ab}})^{G/N}\to G^{\mathrm{ab}}$ (2.2)
is injective. Consider the corestriction and the transfer morphisms
$\mathrm{cor}_{N}:N^{\mathrm{ab}}\to G^{\mathrm{ab}}\qquad\mathrm{transfer}_{N}:G^{\mathrm{ab}}\to N^{\mathrm{ab}}.$
Let $G/N=\cup_{1\leq i\leq[G:N]}a_{i}N$ be a disjoint union of left-cosets with representatives $\{a_{i}\}_{i}$. Then, for any $n\in N$, we have $\mathrm{transfer}_{N}(\mathrm{cor}_{N}(n))=\sum(a_{i}^{-1}na_{i})$ on $N^{\mathrm{ab}}$, i.e.,
$\mathrm{transfer}_{N}\circ\mathrm{cor}_{N}=\sum_{a\in G/N}a\text{-conjugation}$
on $N^{\mathrm{ab}}$. In particular, the restricted morphism $(\mathrm{transfer}_{N}\circ\mathrm{cor}_{N})\mid_{(N^{\mathrm{ab}})^{G/N}}$ coincides with multiplication by $[G:N]$. Since $G$ is ab-torsion-free, $N^{\mathrm{ab}}$ is torsion-free. Hence $\mathrm{transfer}_{N}\circ\mathrm{cor}_{N}$ is injective on $(N^{\mathrm{ab}})^{G/N}$. Therefore, the restricted morphism $(\mathrm{cor}_{N})\mid_{(N^{\mathrm{ab}})^{G/N}}$, which is the morphism 2.2, is also injective. This completes the proof of the claim.
By taking the abelianization of the exact sequence $1\to N\to G\to G/N\to 1$, we have the exact sequence
$(N^{\mathrm{ab}})_{G/N}\to G^{\mathrm{ab}}\to(G/N)^{\mathrm{ab}}\to 1,$
where $(N^{\mathrm{ab}})_{G/N}$ stands for the module of $(G/N)$-coinvariants of $N^{\mathrm{ab}}$. By running over all $N\in\mathcal{N}$, we obtain that the sequence
[…]
is also exact (see [12, Proposition 2.2.4]). Since $G^{\mathrm{ab}}\xrightarrow{\sim}\left(G^{m-1}\right)^{\mathrm{ab}}$, the left-hand morphism of 2.3 is the zero map. The natural morphisms $(N^{\mathrm{ab}})^{G/N}\rightarrow(N^{\mathrm{ab}})_{G/N}\rightarrow G^{\mathrm{ab}}$ induce
[…]
This completes the proof. ∎
@@
Let $G$ be an ab-torsion-free ab-faithful profinite group, and let $m\in\mathbb{Z}_{\geq 2}$. Then both $G$ and $G^{m}$ are center-free.

### v3

@@ 1.3. Proof of the slimness of free $m$-step solvable groups
it suffices to show that $\rho_{H}(g)=1$, i.e., the condition $g\in H^{[m-1]}$ holds, for each such $H$, where $\rho_{H}:H\twoheadrightarrow H^{m-1}$ is the natural surjection. The image of $x$ in the finite quotient $\mathcal{F}^{m}/H$ has finite order. Let $N$ denote this order. Since $g$ commutes with $x^{n}$, it also commutes with $x^{Nn}$, and therefore $\rho_{H}(g)$ commutes with $\rho_{H}(x^{Nn})$. By the Nielsen–Schreier theorem, the inverse image $\tilde{H}$ of $H$ in $\mathcal{F}$ is again a free pro-$\Sigma$ group, and we may choose a free generating set of $\tilde{H}$ that contains $x^{N}$. By Lemma A, we have $H^{m-1}\cong\tilde{H}^{m-1}$. Applying the induction hypothesis for m1m-1 to $\tilde{H}^{m-1}$ and the basis element $x^{N}\in\tilde{H}^{m-1}$, we obtain
$\mathrm{C}_{H^{m-1}}\left((x^{N})^{n}\right)=\overline{\left\langle x^{N}\right\rangle}.$
On the other hand, we have $g\in(\mathcal{F}^{m})^{[m-1]}\subset(\mathcal{F}^{m})^{[2]}\subset H^{[1]}$ and hence $\rho_{H}(g)\in(H^{m-1})^{[1]}$. Note that $\overline{\left\langle x^{N}\right\rangle}$ embeds into $(H^{m-1})^{\mathrm{ab}}$, whereas $(H^{m-1})^{[1]}$ has trivial image there. Therefore,
$\rho_{H}(g)\in\overline{\left\langle x^{N}\right\rangle}\cap(H^{m-1})^{[1]}=1.$
@@
$\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)=\overline{\left\langle x\right\rangle}$
@@
Finally, we consider the case $r=\infty$. Let $J$ be the directed set of finite subsets $X_{j}$ of $X$ such that $x\in X_{j}$. For each $j\in J$, let $\mathcal{F}_{j}$ be the finitely generated free pro-$\Sigma$ group on $X_{j}$ and let $\pi_{j}:\mathcal{F}\twoheadrightarrow\mathcal{F}_{j}$ be the continuous morphism sending generators in $X_{j}$ to themselves and generators in $X\setminus X_{j}$ to the identity element of $\mathcal{F}_{j}$. Additionally, let $\pi_{j}^{(m)}\colon\mathcal{F}^{m}\twoheadrightarrow\mathcal{F}_{j}^{m}$ be the natural projection induced from $\pi_{j}$. Then, by [12, Proposition 3.3.9], we have an isomorphism
@@
Let $g\in\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)$. For each $j$, by the finite-rank case we obtain $\pi_{j}^{(m)}(g)\in\overline{\left\langle\pi^{(m)}_{j}(x)\right\rangle}$. Passing to the inverse limit, we conclude that $g\in\overline{\left\langle x\right\rangle}$. Thus the equality $\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)=\overline{\left\langle x\right\rangle}$ also holds when $r=\infty$. This completes the proof. ∎
We say that a profinite group $G$ is slim if the centralizer $\mathrm{C}_{G}\left(H\right)$ of each open subgroup $H\subset G$ in $G$ is trivial (see [8, Definition 0.1]). We note that slimness implies center-freeness.
@@
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$. Assume $r\neq 1$. Then, for any $m\in\mathbb{Z}_{\geq 2}$, the quotient $\mathcal{F}^{m}$ is slim.
@@
$\mathrm{C}_{\mathcal{F}^{m}}\left(H\right)\subset\mathrm{C}_{\mathcal{F}^{m}}\left(x^{n}\right)\cap\mathrm{C}_{\mathcal{F}^{m}}\left((x^{\prime})^{n^{\prime}}\right)=\overline{\left\langle x\right\rangle}\cap\overline{\left\langle x^{\prime}\right\rangle}=1,$
where the last equality follows from the facts that $\overline{\left\langle x\right\rangle}$ and $\overline{\left\langle x^{\prime}\right\rangle}$ embed into the abelianization $\mathcal{F}^{\mathrm{ab}}$ and are distinct. This completes the proof. ∎
@@ 2. The $m$-step solvable Grothendieck conjecture
In this section, we show that the maximal $m$-step solvable quotients of the geometric étale and tame fundamental groups of hyperbolic curves over a field are center-free (see Theorem 2.9). Moreover, we relate this result to the Grothendieck conjecture. Throughout this section, let $\Sigma$ be a non-empty set of prime numbers. For any profinite group $G$, we write $G^{\Sigma}$ for the maximal pro-$\Sigma$ quotient of $G$.
2.1. Ab-torsion-freeness and ab-faithfulness
@@
In this subsection, we introduce ab-torsion-freeness and ab-faithfulness for profinite groups, and record a strategy for proving center-freeness of maximal $m$-step solvable quotients.
@@
    We say that $G$ is ab-faithful if, for each open subgroup $H$ of $G$ and each open normal subgroup $N$ of $H$, the natural morphism
    $H/N\to\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)$
@@
Remark 2.2.
Let $G$ be a profinite group and let $m\in\mathbb{Z}_{\geq 2}$. For any open subgroup $P$ of $G^{m}$ that contains $(G^{m})^{[m-1]}$, let $\tilde{P}\subset G$ be its inverse image under $G\twoheadrightarrow G^{m}$. Then the natural morphism $\tilde{P}^{\mathrm{ab}}\to P^{\mathrm{ab}}$ is an isomorphism by Lemma A. In particular, the following hold:
  1. (i)
    Assume that $G$ is ab-torsion-free, and let $H$ be an open subgroup of $G^{m}$. If $(G^{m})^{[m-1]}\subset H$, then $H^{\mathrm{ab}}$ is torsion-free.
  2. (ii)
    Assume that $G$ is ab-faithful. Let $H$ be an open subgroup of $G^{m}$ and $N$ an open normal subgroup of $H$. If $(G^{m})^{[m-1]}\subset N$, then the conjugation action of $H/N$ on $N^{\mathrm{ab}}$ is also faithful.
In what follows, we often only need these properties for open subgroups that contain $(G^{m})^{[m-1]}$.
Lemma 2.3.
Let $G$ be an ab-torsion-free profinite group.
@@
    Any closed subgroup of $G$ is ab-torsion-free.
@@
    $G$ is torsion-free.
@@
    For any $m\in\mathbb{Z}_{\geq 1}$, the quotient $G^{m}$ is torsion-free.
  4. (4)
    Let $K$ be a closed normal subgroup of $G$ that is contained in $G^{[1]}$. Then the conjugation action of $G/K$ on $K^{\mathrm{ab}}$ has no nontrivial fixed points.
Proof.
1 Let $K$ be a closed subgroup of $G$. Since $G$ is profinite, we have
$K=\cap_{H}H,$
[…]
2 Let $g\in G$ have finite order. Then the cyclic subgroup $\left\langle g\right\rangle$ is finite and hence closed. By 1, we obtain $\left\langle g\right\rangle^{\mathrm{ab}}=\left\langle g\right\rangle$ is torsion-free, hence $g=1$. Thus $G$ is torsion-free.
3 We prove by induction on $m$. If $m=1$, then the assertion follows from the ab-torsion-freeness of $G$. Assume $m\geq 2$. Then we have the exact sequence:
$1\to(G^{m})^{[m-1]}\to G^{m}\to G^{m-1}\to 1$
By the induction hypothesis, we have that $G^{m-1}$ is torsion-free. Moreover, by 1, the quotient $(G^{m})^{[m-1]}=(G^{[m-1]})^{\mathrm{ab}}$ is also torsion-free. Therefore $G^{m}$ is torsion-free.
4 Let $N$ be an open normal subgroup of $G$. First, we claim that the natural morphism
$(N^{\mathrm{ab}})^{G/N}\to G^{\mathrm{ab}}$ (2.1)
is injective. Indeed, consider the following natural morphism and transfer morphism:
$R_{N}\colon N^{\mathrm{ab}}\to G^{\mathrm{ab}},\qquad\mathrm{transfer}_{N}\colon G^{\mathrm{ab}}\to N^{\mathrm{ab}}$
Let $G/N=\cup_{1\leq i\leq[G:N]}a_{i}N$ be a disjoint union of left cosets with representatives $\{a_{i}\}_{i}$. For each $n\in N$, we have $\mathrm{transfer}_{N}(R_{N}(n))=\sum(a_{i}^{-1}na_{i})$ on $N^{\mathrm{ab}}$, i.e., we have
$\mathrm{transfer}_{N}\circ R_{N}=\sum_{a\in G/N}a\text{-conjugation}$
on $N^{\mathrm{ab}}$. In particular, the restricted morphism $(\mathrm{transfer}_{N}\circ R_{N})\mid_{(N^{\mathrm{ab}})^{G/N}}$ coincides with multiplication by $[G:N]$. Since $G$ is ab-torsion-free, the group $N^{\mathrm{ab}}$ is torsion-free. Hence $\mathrm{transfer}_{N}\circ R_{N}$ is injective on $(N^{\mathrm{ab}})^{G/N}$. Therefore, the restricted morphism $(R_{N})\mid_{(N^{\mathrm{ab}})^{G/N}}$, which is the morphism (2.1), is also injective. This completes the proof of the claim.
Let $\mathcal{N}$ be a set of all open subgroups that contain $K$. By running over all $N\in\mathcal{N}$, we have
[…]
Therefore, this claim implies that the natural morphism
$(K^{\mathrm{ab}})^{G/K}\to G^{\mathrm{ab}}$
is also injective.
By taking the abelianization of the exact sequence $1\to K\to G\to G/K\to 1$, we have the exact sequence
$(K^{\mathrm{ab}})_{G/K}\to G^{\mathrm{ab}}\to(G/K)^{\mathrm{ab}}\to 1,$ (2.2)
where $(K^{\mathrm{ab}})_{G/K}$ stands for the module of $G/K$-coinvariants of $K^{\mathrm{ab}}$. Since $K\subset G^{[1]}$, the natural morphism $G^{\mathrm{ab}}\xrightarrow{\sim}(G/K)^{\mathrm{ab}}$ is an isomorphism. Hence the left-hand morphism of (2.2) is the zero map. The above claim implies that the composition of these natural morphisms
$(K^{\mathrm{ab}})^{G/K}\hookrightarrow K^{\mathrm{ab}}\twoheadrightarrow(K^{\mathrm{ab}})_{G/K}\overset{0}{\to}G^{\mathrm{ab}}.$
is injective. Therefore, we obtain $(K^{\mathrm{ab}})^{G/K}=1$. This completes the proof. ∎
@@
Let $G$ be an ab-faithful profinite group.
  1. (1)
    $G$ is center-free.
  2. (2)
    For any $m\in\mathbb{Z}_{\geq 1}$, we have $\mathrm{Z}\left(G^{m}\right)\subset(G^{m})^{[m-1]}$.
@@
1 Let $N$ be an open normal subgroup of $G$. Then we have
@@
$\mathrm{Z}\left(G\right)\subset\ker\Bigl(G/N\longrightarrow\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)\Bigr).$
By ab-faithfulness (applied to the pair $(H,N)=(G,N)$), the above morphism is injective, hence the kernel is trivial. Therefore, by running over all such $N$, we obtain
$\mathrm{Z}\left(G\right)\subset\bigcap_{N}N=\{1\}.$
Thus the group $G$ is center-free.
2 The proof is essentially the same as the proof of 1. Let $N$ be an open normal subgroup of $G^{m}$ that contains $(G^{m})^{[m-1]}$. Then, by Remark 2.2ii, we have $\mathrm{Z}\left(G^{m}\right)\subset\ker\Bigl(G^{m}/N\longrightarrow\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)\Bigr)$. By ab-faithfulness (applied to the pair $(H,N)=(G^{m},N)$), this morphism is injective, hence the kernel is trivial. Therefore, by running over all such $N$, we obtain
$\mathrm{Z}\left(G^{m}\right)\subset\bigcap_{N}N=(G^{m})^{[m-1]}.$
The following proposition is the main result of this subsection.
Proposition 2.5.
Let $G$ be an ab-torsion-free ab-faithful profinite group. Then, for any $m\in\mathbb{Z}_{\geq 2}$, the quotient $G^{m}$ is torsion-free and center-free.
Proof.
The torsion-freeness follows from Lemma 2.33. We show the center-freeness. For any $a\in G^{m}$, the condition $a\in\mathrm{Z}\left(G^{m}\right)$ is equivalent to the condition that $gag^{-1}=a$ for every $g\in G^{m}$, and hence
$\mathrm{Z}\left(G^{m}\right)\cap(G^{m})^{[m-1]}=((G^{m})^{[m-1]})^{G^{m-1}}.$
By applying Lemma 2.34 to $K=G^{[m-1]}$, the right-hand side is trivial. On the other hand, by Lemma 2.42, we have
$\mathrm{Z}\left(G^{m}\right)\subset(G^{m})^{[m-1]}.$
Thus $G^{m}$ is center-free. ∎
2.1.2.
As we have already seen, a free pro-$\Sigma$ group is slim. Since slimness is stronger than center-freeness, it is natural to ask whether it also holds for the maximal $m$-step solvable quotients of ab-faithful and ab-torsion-free profinite groups. At the time of writing, the author does not know whether such groups are slim in general. However, the following fact can be proved:
Proposition 2.6.
Let $G$ be an ab-torsion-free and ab-faithful profinite group. Then, for any $m\in\mathbb{Z}_{\geq 1}$ and any open subgroup $H$ of $G^{m}$, we have
$\mathrm{C}_{G^{m}}\left(H\right)\subset(G^{m})^{[m-1]}.$
Proof.
Let $H\subset G^{m}$ be an open subgroup, and take $c\in\mathrm{C}_{G^{m}}\left(H\right)$. Let $N\trianglelefteq_{\mathrm{open}}G^{m}$ be an open normal subgroup such that $(G^{m})^{[m-1]}\subset N$. Let $H_{N}\subset N^{\mathrm{ab}}$ be the image of $H\cap N$ under the natural morphism $N\twoheadrightarrow N^{\mathrm{ab}}$. Since $c$ centralizes $H$, conjugation by $c$ is trivial on $H\cap N$, hence it is also trivial on $H_{N}$. As $H\cap N$ is open in $N$, the subgroup $H_{N}$ is open (equivalently, of finite index) in $N^{\mathrm{ab}}$. Therefore $N^{\mathrm{ab}}/H_{N}$ is a torsion group, and hence the natural morphism
$H_{N}\otimes_{\mathbb{Z}}\mathbb{Q}\xrightarrow{\sim}N^{\mathrm{ab}}\otimes_{\mathbb{Z}}\mathbb{Q}$
is an isomorphism. It follows that conjugation by $c$ acts trivially on $N^{\mathrm{ab}}\otimes_{\mathbb{Z}}\mathbb{Q}$. By Remark 2.2i, the group $N^{\mathrm{ab}}$ is torsion-free. Thus the natural morphism $N^{\mathrm{ab}}\hookrightarrow N^{\mathrm{ab}}\otimes_{\mathbb{Z}}\mathbb{Q}$ is injective. Consequently, conjugation by $c$ is already trivial on $N^{\mathrm{ab}}$. On the other hand, by Remark 2.2ii, the conjugation action of $G^{m}/N$ on $N^{\mathrm{ab}}$ is faithful, i.e., we have
$\ker\Bigl(G^{m}\to\operatorname{Aut}(N^{\mathrm{ab}})\Bigr)=N.$
Since $c$ acts trivially on $N^{\mathrm{ab}}$, we obtain $c\in N$. By running over all such $N$, we conclude that
$c\in\bigcap_{N}N=(G^{m})^{[m-1]}.$
This completes the proof. ∎
2.2. Proof of the center-freeness of the maximal $m$-step solvable quotients of the geometric fundamental groups of hyperbolic curves
2.2.1.
In this subsection, we show that the maximal $m$-step solvable quotients of the geometric fundamental groups of smooth curves are center-free. We always assume that smooth curves are geometrically connected. For any smooth curve $X$ over a field $k$, we write
$\pi_{1}^{\mathrm{\acute{e}t}}(X,\ast)\qquad(\text{resp. }\pi_{1}^{\mathrm{tame}}(X,\ast))$
for the étale fundamental group (resp. tame fundamental group) of $X$, where $\ast:\operatorname{Spec}(\Omega)\to X$ denotes a geometric point of $X$ and $\Omega$ denotes an algebraically closed field. The fundamental group depends on the choice of base point only up to inner automorphisms, and therefore we omit the choice of base point below.
2.2.2.
We define the Euler characteristic of $X$ by
$\chi(X)\coloneq\sum_{i=0}^{2}(-1)^{i}\dim_{\mathbb{Q}_{\ell}}(\mathrm{H}_{\mathrm{\acute{e}t}}^{i}(X,\mathbb{Q}_{\ell}))=\begin{cases}2-\dim_{\mathbb{Q}_{\ell}}(\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(X,\mathbb{Q}_{\ell}))&(X:\ \text{proper}),\\ 1-\dim_{\mathbb{Q}_{\ell}}(\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(X,\mathbb{Q}_{\ell}))&(X:\ \text{affine}).\end{cases}$ (2.3)
If $X$ is of type $(g,r)$, then a straightforward calculation shows that
$\chi(X)=2-2g-r.$
We say that $X$ is hyperbolic if $\chi(X)<0$ (equivalently, if $(g,r)\notin\{(0,0),(0,1),(0,2),(1,0)\}$). The basic fact about hyperbolicity is that, if $\ell$ is a prime number different from the characteristic of $k$, then
[…]
(see [14, Corollary 1.4]).
@@
Let $k$ be a separably closed field and let $\ell$ be a prime number different from the characteristic of $k$. Let $X$ be a hyperbolic curve over $k$, and let $f\colon Y\to X$ be a finite étale Galois covering with Galois group $\Gamma\coloneq\operatorname{Gal}(Y/X)$. Then the natural action
$\Gamma\curvearrowright\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})$
is faithful.
@@
Replacing $k$ by an algebraic closure does not change the statement. Hence we may assume that $k$ is algebraically closed. Let
$\Gamma_{0}\coloneq\ker\!\Bigl(\Gamma\longrightarrow\operatorname{Aut}_{\mathbb{Q}_{\ell}}\bigl(\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})\bigr)\Bigr),$
and put $Y_{0}\coloneq Y/\Gamma_{0}$. Then $Y\to Y_{0}$ is a finite étale Galois covering with Galois group $\Gamma_{0}$. Consider the Hochschild–Serre spectral sequence for the Galois covering $Y\to Y_{0}$ with coefficients $\mathbb{Q}_{\ell}$:
$E_{2}^{p,q}=\mathrm{H}^{p}\!\bigl(\Gamma_{0},\,\mathrm{H}_{\mathrm{\acute{e}t}}^{q}(Y,\mathbb{Q}_{\ell})\bigr)\ \Longrightarrow\ \mathrm{H}_{\mathrm{\acute{e}t}}^{p+q}(Y_{0},\mathbb{Q}_{\ell}).$
Then we obtain the associated five-term exact sequence
$0\to\mathrm{H}^{1}(\Gamma_{0},\mathbb{Q}_{\ell})\to\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y_{0},\mathbb{Q}_{\ell})\to\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})^{\Gamma_{0}}\to\mathrm{H}^{2}(\Gamma_{0},\mathbb{Q}_{\ell}).$
Here $\Gamma_{0}$ is finite and hence $\mathrm{H}^{1}(\Gamma_{0},\mathbb{Q}_{\ell})=\mathrm{H}^{2}(\Gamma_{0},\mathbb{Q}_{\ell})=0$. Therefore, the restriction morphism
$\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y_{0},\mathbb{Q}_{\ell})\xrightarrow{\ \sim\ }\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})^{\Gamma_{0}}$
is an isomorphism. By definition of $\Gamma_{0}$, the $\Gamma_{0}$-action on $\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})$ is trivial, hence
$\dim_{\mathbb{Q}_{\ell}}(\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y_{0},\mathbb{Q}_{\ell}))=\dim_{\mathbb{Q}_{\ell}}(\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})).$
Since the morphism $Y\to Y_{0}$ is finite, the curve $Y$ is proper if and only if $Y_{0}$ is proper. Hence (2.3) implies that $\chi(Y_{0})=\chi(Y)$.
On the other hand, by the Riemann–Hurwitz theorem for their compactifications of $Y\to Y_{0}$ (see [5, Corollary 2.4]), we have the inequality
$\chi(Y)\leq d\cdot\chi(Y_{0}),$
where $d\coloneq\#\Gamma_{0}$. By the hypothesis, the curve $X$ is hyperbolic, and hence $\chi(Y)<0$. This implies that $d=1$, i.e., the action $\Gamma\curvearrowright\mathrm{H}_{\mathrm{\acute{e}t}}^{1}(Y,\mathbb{Q}_{\ell})$ is faithful. ∎
@@
Let $X$ be a hyperbolic curve over a field $k$. Assume that $k$ is a separably closed field and that $\Sigma$ contains a prime number different from the characteristic of $k$. Then the groups $\pi_{1}^{\mathrm{\acute{e}t}}(X)^{\Sigma}$ and $\pi_{1}^{\mathrm{tame}}(X)^{\Sigma}$ are both ab-torsion-free and ab-faithful.